# SPrimary Studies

### v1

@@ 4.4 Phase III: Data Extraction
Our data extraction phase targeted certain relevant information, listed in Table 3. In addition to the metadata about the paper (prefixed by C0 in the table), e.g., its title and the authors name, we extracted the following information: the GDPR concepts and principles covered in the primary studies (C1), the traced personal data (C2), the targeted app store(s) and operating system(s) (C3), the goals and contributions of the primary studies (C4), the applied data analysis methods (C5), the proposed solutions and their enabling technologies (C6), the data used (C7), and whether any artifact was made publicly available (C8). We extracted this information by thoroughly reviewing the 60 papers marked as relevant in the previous phase. Our data extraction process involved two steps, described below.
@@
C1.1 GDPR Principles Key GDPR principles mentioned in the paper
C1.2 Privacy Concepts Key GDPR privacy concepts mentioned in the paper according to our reference model §
@@
C6.3 Program Analysis Whether the paper employs program analysis techniques.
C6.3.1 Method The concrete program analysis method(s) applied in the paper.
C6.4 Outcome The final outcome of the paper.
@@
    See Section 2.
@@ 4.4.1 Preliminary Information Extraction
In this step, we primarily reviewed the abstract and the first sections (usually called “introduction”, “background”, and “related work”). The purpose of this step is to extract generic relevant information related to the main focus of the primary study. Concretely, we extracted the main goal, the contributions, the research questions, the outcomes, and information on online availability of resources, if applicable. We also identified the covered GDPR principles and privacy concepts according to our reference model. We noted that some papers lacked details concerning, e.g., contributions or research questions. When such details were not clearly stated in the analyzed papers, we labeled the respective information as “not stated”.
@@ 5 Findings
5.1 RQ1. How well does the literature on privacy in mobile apps cover the GDPR privacy metadata types?
To answer RQ1, we mapped the privacy concepts highlighted in the primary studies to the information types in our reference model (see Section 2). This mapping activity involved labeling the primary study with the specific information types, if the study mentions or discusses that information type in any way. Multiple information types can obviously be assigned to a single analyzed study, i.e., the mapping is a multi-labeling task. The mapping was primarily conducted by the first author. To ensure the quality of the resulting labeling, the second and third authors independently validated a subset of 10% primary studies each. We then measured the agreement among the information types identified by the first author and the types identified by the second or the third author using the Krippendorff’s alpha coefficient [41] over this subset. We obtained an agreement of 0.94, indicating a high agreement.
@@
According to our results, 49 out of 60 papers ($\approx$82%) mention one or more GDPR privacy concepts and 25 papers ($\approx$42%) mention one or more GDPR principles. 51 out of 60 papers ($\approx$85%) have been published after 2020, with half of them published in 2023. These percentages indicate the increase of interest in investigating the GDPR privacy concepts in various SE contexts, as we will show later in this section. Table 4 presents the information types corresponding to the GDPR concepts, alongside the analyzed papers discussing them. Note that the total number of papers exceeds 49, since the same paper could discuss multiple information types. Fig. 3 demonstrates the attention given to these information types by showing the number of papers that mention them.
In our analysis, we observed that some primary studies investigate generic GDPR principles. For example, the data minimisation principle refers to collecting only the data necessary to run a service (Article 5(1)(c) of GDPR). We observe that the SE literature often discusses the following GDPR principles: (1) lawfulness, fairness, and transparency; (2) purpose limitation; (3) data minimisation; (4) accuracy; (5) storage limitation; (6) integrity and confidentiality; and (7) accountability [42, 43]. To this end, we note that the reference model we use to do this review does not explicitly refer to GDPR principles. Instead, the model focuses on obligations and rights stipulated in the GDPR that are directly relevant to compliance. We still decided to report on GDPR principles in addition to the information types identified according to the reference model. Our findings are therefore informative to the research community also with respect to GDPR principles in mobile apps literature.
Table 4: Mentions of Information Types in the Literature (RQ1)
Metadata Level§ Total Papers
[…]
    Level corresponds to the hierarchy in the reference model as shown in Fig. 1: L1 is the generic metadata (shaded teal-blue), while L2 (shaded light green) and L3 (shaded gray) provide specializations to L1.
Figure 3: Coverage of GDPR Privacy Concerns in the Literature (RQ1).
We observe that the literature is dominated by three information types. The most investigated information type (in 31 primary studies) is the PERSONAL DATA ORIGIN, which refers to the personal data collection mechanisms, e.g., whether the personal data is collected through cookies. The second is RECIPIENTS, which refers to the sharing of personal data with entities or organizations other than the controller, e.g., with third parties. Finally, LEGAL BASIS is another well-investigated information type, where the literature exclusively focus on CONSENT.
We note that, while CONSENT has a significant value since it authorizes a software system (or mobile app in this case) to process personal data of individuals, other legal bases such as LEGITIMATE INTEREST can impact the implementation of certain actions related to users’ data. For instance, the deletion of personal data upon the users’ request can be postponed or declined because of the LEGAL OBLIGATION based on which personal data was collected in the first place. Personal data can further be processed based on lawful PUBLIC INTEREST (e.g., public health or historical research purposes).
Another key observation is that, despite the significant attention given to CONSENT—since it concerns individuals (i.e., users of a given mobile app)— very little attention has been given to DATA SUBJECT RIGHT, concerning individuals’ rights over their personal data, another user-centered information type that is highly-relevant for the development of legally-compliant mobile apps. Examples of such rights include WITHDRAW CONSENT to enable the user to revoke their consent given in the first place, the right to ACCESS where users should be able to access their personal data collected over a certain period of time, or the right to ERASURE enabling users to request the deletion of their personal data. Allowing individuals to practice their rights is fundamental according to GDPR. If such legal requirements are well-captured and appropriately implemented in mobile apps, individuals can easily practice their rights through, e.g., direct interaction within the app.
Considering that the most well-covered information type is PERSONAL DATA ORIGIN, we looked closely at the types of personal data that are being discussed in the primary studies. Broadly, there are three categories of personal data types. The first category concerns data about individuals, which are not app-dependent, i.e., they do not change depending on which app is being used. Examples in this category include fingerprints, name, phone number, home address, gender, age, bank details, health information, photos, and salary. The data in this case are often collected directly from the user of an app, corresponding therefore to DIRECT information type. The second category concerns data about the user of the mobile app, e.g., user name, password, phone calls, SMS, contacts. Such data often varies across different apps, depending on, for example, the domain and the purpose of the app. The last category concerns data related to the mobile device, e.g., geolocation information, IP address, connectivity status, SIM card ID, addresses of nearby hotspots. We believe it is important to make this distinction at the time of developing the app, to have a better understanding of what data will be collected and hence what requirements are necessary to fulfill in order to develop apps that comply with GDPR. We remind the reader that the GDPR imposes obligations for protecting personal data. It requires more strict obligations (e.g., security mechanisms) when sensitive personal data (e.g., health information) is collected compared to less or non sensitive data, such as the user name invented by an individual for using a specific mobile app.
@@
Requirements pertaining to some information types are likely not implementable, i.e., no software component is expected to deal with them. Such information types are expected to receive very little attention in the SE literature, due to the lack of software relevance. However, these information types often entail legal requirements affecting the involved parties in the data collection and processing activities, such as the mobile software development companies. To be fully GDPR compliant, these requirements must be comprehensively specified in legal agreements such as privacy policies or data processing agreements established among the involved parties. Examples of such information types include PD PROVISION OBLIGED, referring to the failure of providing some services as a consequence for not providing certain personal data to the controller. Another example is CONTROLLER REPRESENTATIVE, which defines a representative entity of the controller when the controller is located outside the EU.
@@
RQ2 investigates the main research problems related to GDPR compliance addressed by the SE community. The main objectives observed in our work can be summarized across eight themes, discussed below. Table 5 lists the primary studies under each theme.
Table 5: Overview of Goals Observed in the Literature (RQ2)
Theme Total Papers
[…]
5.2.1 T1. Tracing data through SOURCE and SINK methods
This theme is frequent in SE, as primary studies often trace certain personal data (SOURCE) to identify potential data leaks. Identifying which personal data to trace is a relevant research topic.
@@
Fostering this research theme is essential for understanding whether the mobile app is GDPR compliant when collecting personal data from the user. Existing work under this theme focuses on analyzing source or byte code of mobile apps and does not necessarily investigate privacy policies.
@@
We believe that this theme crosses multiple information types, including PERSONAL DATA CATEGORY, PD ORIGIN, and CONSENT. Sharing data with third-party libraries is a closely related theme; in our reference model, this theme concerns the information type RECIPIENT. In brief, the GDPR requires all recipients to be listed in the privacy policies of mobile apps. If the app is sharing personal data with third-party libraries, then this has to be fully transparent.
5.2.2 T2. App permission requests
@@
This theme investigates the access to personal data granted to the mobile app through permission requests. As mobile apps are becoming more pervasive, this theme is important for understanding what personal information the mobile app can access or collect indirectly through permission requests, e.g., to access information in other apps. Typical example is when the user ought to share location information to be able to use various apps such as navigation or fitness apps. This theme is relevant to the information types: PERSONAL DATA SECURITY, RECIPIENTS and PD ORIGIN.
5.2.3 T3. Security and privacy of minors.
@@
This theme focuses on minors, with an emphasis on the threats, risks, and consent needed to ensure that their rights are adequately protected. This theme maps directly to the CHILDREN information type.
5.2.4 T4. Informed consent
@@
Consent is a well-studied concept. The GDPR emphasizes on collecting explicit consent from individuals, enabling them to make informed decisions with regard to the handling of their personal data. However, according to the reference model we apply in this SLR, CONSENT is only one of the several possible legal bases for data collection and processing.
@@
CONSENT is also the easiest to verify in software as it should be translated into interaction between the mobile app and the user. To investigate whether the practices in the mobile app code are GDPR compliant with respect to some other legal bases, a more comprehensive analysis of the app context must be conducted; such an analysis would not be necessarily limited to, e.g., the app’s source code. We note that, despite not being popular in the SE literature, legal bases other than consent can play an important role in mobile apps development decisions.
5.2.5 T5. Security/privacy awareness
@@
This theme surveys the users of mobile apps to study how well individuals are aware of their rights as well as common practices concerning privacy and security issues, often perceived in mobile apps’ behaviors. While this theme highlights the importance of individuals’ awareness about data protection, it cannot be directly mapped to our reference model.
@@
5.2.6 T6. Consistency between apps and their privacy policies
@@
Some papers investigate the consistency between the content of the privacy policy associated with an app against the actual behavior of that app. However, this research theme is often scoped to a use case such as the collection and sharing of sensitive personal data with third parties.
@@
Advancing this theme requires in-depth analysis of at least the following information types: PERSONAL DATA CATEGORY, LEGAL BASIS, RECIPIENTS, PD ORIGIN, and DATA SUBJECT RIGHT. These information types facilitate the understanding of what personal data is collected, the source where it is obtained from, with whom it will be shared, what rights individuals can have on their personal data, and how these rights are implemented.
5.2.7 T7. Detecting violations of data minimization principle
Data minimization is an important guiding principle in GDPR: according to this principle, an app should not collect more personal data than necessary for its use.
@@
To ensure that no personal data is unnecessarily collected, the findings related to these two information types should then be cross-checked against PERSONAL DATA CATEGORY, indicating the personal data types collected, and PD ORIGIN, indicating the source where the data is collected from.
@@
This theme focuses on making privacy policies more accessible to individuals by providing informative summaries or categorization of the policies.
While the primary studies under this theme cannot be one-to-one mapped to information types in our reference model, we believe that the reference model can be useful to advance this research theme as it provides a structured hierarchy of the necessary details that must be identified in privacy policies.
@@
RQ3 explores the various contributions identified in the literature concerning GDPR. We distinguish in this SLR between technical and non-technical contributions: the former refers to proposing automated solutions for addressing GDPR privacy concerns, whereas the latter are concerned with sharing insights or recommendations derived from, e.g., surveys, aiming at advancing the research knowledge.
@@
Most of these studies are qualitative, relying on manual means for analyzing certain privacy concerns. However, some studies used simple tool support such as keyword-based or string search for extracting indicator terms, or applied third-party tools such as Mobile Security Framework (MobSF) [^9] https://github.com/MobSF/Mobile-Security-Framework-MobSF, FlowDroid [^10] https://github.com/secure-software-engineering/FlowDroid or Wireshark [^11] https://www.wireshark.org/ for static or dynamic code analysis. Such an automation was strictly used to support their empirical studies or surveys.
[…]
@@
The remaining 25 primary studies make technical contributions. They proposed automation strategies by leveraging different enabling technologies: model-driven engineering was used to enhancing business process modeling notation with privacy considerations [S6], static and dynamic program analysis were used for detecting violations or vulnerabilities in mobile code [S35, S41, S46, S51, S59], NLP was employed for analyzing privacy policies and app reviews [S36, S44, S53, S57], machine learning (ML) and deep learning (DL) were used for classifying privacy policy provisions [S30] and detecting source and sink method detection [S1]. About half of the primary studies under this category leveraged a combination of these enabling technologies, particularly NLP, ML, DL, and static or dynamic program analysis. These studies pursued the following objectives:
  • identifying data recipients [S8],
  • detecting privacy issues in app reviews [S16],
  • detecting sensitive information disclosure in mobile apps [S20],

### v2

@@ 4.4 Phase III: Data Extraction
Our data extraction phase targeted certain relevant information, listed in Table 5. In addition to the metadata about the paper (prefixed by C0 in the table), e.g., its title and the authors name, we extracted the following information: the GDPR concepts and principles covered in the primary studies (C1), the traced personal data (C2), the targeted app store(s) and operating system(s) (C3), the goals and contributions of the primary studies (C4), the applied data analysis methods (C5), the proposed solutions and their enabling technologies (C6), the data used (C7), and whether any artifact was made publicly available (C8). We extracted this information by thoroughly reviewing the 60 papers marked as relevant in the previous phase. Our data extraction followed an integrated coding approach [48], combining both inductive and deductive analysis. We started our analysis with a predefined set of codes derived from the concepts outlined in our reference model. However, as not all primary studies explicitly stated these concepts explicitly, we also based our extraction on interpretive analysis to identify implicit evidence pertinent to the concepts. Consequently, our findings draw both on the verbatim content of the studies and on our interpretation in relation to the reference model. This hybrid method enabled a more comprehensive and grounded analysis of the data.
The fields prefixed with C1.* are intended for addressing RQ1, while those prefixed with C2.*, C3.*, and C4.* are intended for addressing RQ2, with C3 providing a focused view on the possible app market restrictions scoped in the primary studies. Fields prefixed with C4.*, C5.*, C6.* are designed to answer RQ3, while those prefixed with C7.* and C8.* are intended to answer RQ4.
Regarding C1 fields, our coding activity involved labeling the primary studies with the relevant set of GDPR principles listed in Art. 5 (C1.1), the set of concepts from our reference model (C1.2), and any other GDPR-relevant concern not captured by the first two fields (C1.3). The motivation for including C1.3 is to make our findings more informative to the research community. A single primary study can obviously contain mentions of multiple concepts, principles, other concerns, or a combination thereof. The remaining fields capture information that are not strictly dependent on GDPR principles or concepts.
The coding was primarily performed by the first author. To ensure the reliability of the resulting labels, the second and third authors independently validated two disjoint subsets, each covering 10% primary studies. We then measured the agreement between the concepts identified by the first author and those identified by the second and third authors using the Krippendorff’s alpha coefficient [49]. We obtained an agreement value of 0.94, indicating high agreement among the coders.
@@
C1.1 GDPR Principles Key GDPR principles (according to GDPR Art. 5) mentioned in the paper.
C1.2 Privacy Concepts Key GDPR privacy concepts mentioned in the paper according to our reference model§.
C1.3 Other Concerns Other GDPR-related concerns mentioned in the paper.
@@
Our data extraction process involved two steps, described below.
@@ 4.4.1 Preliminary Information Extraction
In this step, we primarily reviewed the abstract and the “introduction” section (together with the “background” and “related work/state of the art” sections if needed to better understand the paper). The purpose of this step is to extract generic relevant information related to the main focus of the primary study. Concretely, we extracted the main goal, the contributions, the research questions, and information on online availability of resources, if applicable. We also identified the covered GDPR principles and privacy concepts. When such details, e.g., concerning contributions or research questions, were not clearly stated in the analyzed papers, we labeled the respective information as “not stated”.
@@ 4.4.2 Detailed Information Extraction
4.5 Data Availability
We make the raw data of our study available to reviewers in an annex [29]. We plan to make it publicly available upon acceptance of the paper.
@@ 5 Findings
5.1 RQ1. How well does the literature on privacy in mobile apps cover the GDPR privacy concepts?
To answer RQ1, we mapped the privacy concerns highlighted in the primary studies to the concepts in our reference model and the list of principles from Art. 5 of GDPR (see Section 2).
According to our observations, 48 out of 60 papers (80%) mention one or more GDPR privacy concepts, 19 papers ($\approx$32%) mention one or more principles (from Art. 5), and 18 papers (30%) mention other privacy concerns, besides the concepts and principles. Other concerns are mostly related to privacy or security in a broader sense than what is in the GDPR, including privacy by design, data anonymization, data transfer to third parties [^10] Note that “Third Party” in Table 6 indicate collecting data from third-parties (not transferring to them)., and notification in case of data breach. Finally, five papers ($\approx$8.3%) lack mentions of any GDPR principle or concept, but they refer to privacy at large, data protection at large, privacy protection, privacy by design, security, and security by design. These papers are typically about privacy risks rather than rights and obligations under GDPR.
@@
51 out of 60 papers ($\approx$85%) have been published after 2020, with half of them published in 2023. These percentages indicate the increase of interest in investigating the GDPR privacy concepts in various SE contexts, as we will show later in this section. Table 6 presents the mapping between the GDPR concepts and the analyzed papers. Following our reference model, we organize the concepts in the table into three layers (L1, L2, and L3), with L3 being a specialization of L2, which is in turn a specialization of L1. We use separators between the different L1-concepts (grouped with their L2 and L3 specializations). Following this, we count the number of primary studies as follows: whenever a concept is identified at levels L2 or L3, we also count the paper in the parent concept (L1 or L2, respectively), e.g., a paper discussing WITHDRAW CONSENT (L2) is by default also discussing DATA SUBJECT RIGHT (L1), since the former is a specialization of the latter. Note that the total number of papers exceeds 48, since the same paper could discuss multiple concepts.
Table 6: Mentions of GDPR Concepts in the Literature (RQ1)
Concept Level§ Total Papers
Recipients L1 26 [2, 3, 7, 8, 12, 13, 14, 17, 23, 28, 29, 31, 35, 38, 39, 40, 41, 44, 46, 47, 48, 51, 52, 53, 54, 59]
Children L1 7 [2, 9, 10, 24, 32, 37, 40]
PD Security L1 3 [2, 11, 40]
PD Time Stored L1 9 [2, 7, 11, 14, 23, 30, 37, 40, 53]
Auto Decision Making L1 3 [12, 14, 44]
Processing Purposes L1 9 [2, 7, 11, 14, 17, 23, 30, 37, 43]
DPO L1 1 [30]
Contact L2 1 [30]
Controller L1 2 [30, 40]
Contact L2 2 [30, 40]
Controller Representative L1 0
Data Subject Right L1 9 [2, 10, 14, 30, 37, 40, 44, 53, 56]
Withdraw Consent L2 3 [44, 53, 56]
PD Category L1 13 [11, 14, 15, 41, 48, 51, 52, 54, 55, 56, 57, 58, 60]
Special L2 3 [14, 41, 58]
Type L2 1 [54]
PD Origin L1 32 [2, 3, 7, 11, 12, 13, 14, 15, 17, 23, 28, 29, 30, 31, 34, 35, 37, 38, 39, 40, 41, 43, 44, 48, 51, 52, 53, 54, 56, 58, 59, 60]
Direct L2 31 [2, 3, 7, 11, 12, 13, 14, 15, 17, 23, 28, 29, 30, 31, 34, 35, 37, 38, 39, 40, 41, 43, 48, 51, 52, 53, 54, 56, 58, 59, 60]
Indirect L2 3 [44, 53, 54]
Third Party L3 2 [53, 54]
Cookie L3 1 [44]
Transfer outside Europe L1 1 [53]
Safeguards L2 1 [53]
Legal Basis L1 21 [1, 5, 7, 11, 14, 22, 32, 34, 38, 39, 41, 42, 43, 44, 45, 48, 49, 50, 53, 56, 58]
Consent L2 21 [1, 5, 7, 11, 14, 22, 32, 34, 38, 39, 41, 42, 43, 44, 45, 48, 49, 50, 53, 56, 58]
Legitimate Interest L2 1 [49]
Vital Interest L2 1 [49]
Contract L2 1 [49]
Public Function L2 1 [49]
Legal Obligation L2 1 [49]
PD Provision Obliged L1 0
    Level corresponds to the hierarchy in the reference model as shown in Fig. 1.
@@
We observed from the table that the only L1 concepts which exist in the reference model, yet were not identified in the literature are CONTROLLER REPRESENTATIVE and PD PROVISION OBLIGED, as shown in Table 6. The former refers to an administrative representative acting on behalf of the data controller, while the latter concerns the denial of services when certain personal data are not shared with the controller. From a research standpoint, these concepts seem not to be of particular relevance. A similar observation can be made with regards to missing L2 and L3 privacy concepts (that are not listed in the table due to space limitations). These concepts are largely associated with requirements that are unlikely to be implementable; no software component is expected to address them and hence they are likely to receive little attention from the SE community. Nonetheless, such concepts typically entail legal obligations affecting the involved parties in the data collection and processing activities, including the mobile software development companies. To be fully GDPR compliant, these concepts must be comprehensively specified in legal agreements, such as privacy policies or data processing agreements, established among the involved parties.
We observe that the literature is dominated by three privacy concepts. The most investigated concept (in 32 primary studies) is the PERSONAL DATA ORIGIN (L1), which refers to the source for the personal data collection, e.g., either directly provided by the data subject (L2) or indirectly (L2), for instance, through cookies (L3). The second is RECIPIENTS (L1) (in 26 studies), which refers to the sharing of personal data with entities or organizations other than the controller, e.g., with third parties. Finally, LEGAL BASIS (L1) is another well-investigated concept (in 21 studies), where the literature exclusively focus on CONSENT (L2): all of the 21 studies focus on CONSENT (L2); only one primary study had mentions of the remaining LEGAL BASES (L1).
We note that, while CONSENT has a significant value since it authorizes a software system (or mobile app in this case) to process personal data of individuals, other legal bases such as LEGITIMATE INTEREST (L2) can impact the implementation of certain actions related to users’ data. For instance, the deletion of personal data upon the users’ request can be postponed or declined because of the LEGAL OBLIGATION (L2) based on which personal data was collected in the first place. Personal data can further be processed based on lawful PUBLIC INTEREST (L2) (e.g., public health or historical research purposes).
Another key observation is that, despite the significant attention given to CONSENT(L2)—since it concerns individuals (i.e., users of a given mobile app)—very little attention has been given to (DATA SUBJECT RIGHT (L1) (only in nine primary studies, with three specifically on WITHDRAW CONSENT (L2)). DATA SUBJECT RIGHTS (L1) concern individuals’ rights over their personal data, another user-centered concept that is highly-relevant for the development of legally-compliant mobile apps. The primary studies remained at a generic level (L1) except for WITHDRAW CONSENT (L2) (explicitly mentioned in three papers) to enable the user to revoke their consent given in the first place. Essential rights such as the right to ACCESS (L2) where users should be able to access their personal data collected over a certain period of time, or the right to ERASURE (L2) enabling users to request the deletion of their personal data were not specifically considered. Allowing individuals to practice their rights is fundamental according to GDPR. If such legal requirements are well-captured and appropriately implemented in mobile apps, individuals can easily practice their rights through, e.g., direct interaction within the app.
Overall, we observe that, apart from CONSENT (L2) and WITHDRAW CONSENT (L2), the majority of the studies discuss GDPR concepts at a high level (mostly L1), overlooking specific details (L2 and L3). We note that a privacy concern being well-covered in the literature does not necessarily imply that it has been resolved. Rather, it reflects the growing interest in and awareness of the community of this concern. Conversely, under-explored concerns, or the lack of in-depth analysis, may indicate limited awareness, which could, in turn, lead to non-compliance practices.
Considering that the most well-covered concept is PERSONAL DATA ORIGIN (L1), and PERSONAL DATA CATEGORY (L1) also being well represented (13), we looked closely at the types of personal data (C2.1 in Table 5) that are being discussed in the primary studies. Broadly, there are three categories of personal data types (related to PERSONAL DATA CATEGORY (L1). The first category concerns data about individuals, which are not app-dependent, i.e., they do not change depending on which app is being used. Examples in this category include name, phone number, home address, gender, age, bank details, health information, photos, and salary. The data in this case are often linked to the DIRECT (L2) concept, i.e., personal information collected directly from the user of an app, (mentioned in 30 primary studies out of the 32 discussing PERSONAL DATA ORIGIN (L1)). The second category concerns data about the user of the mobile app, e.g., user name, password, phone calls, SMS, contacts. Such data often varies across different apps, depending on, for example, the domain and the purpose of the app. The last category concerns data related to the mobile device, e.g., geolocation information, IP address, connectivity status, SIM card ID, addresses of nearby hotspots. We believe it is important to make this distinction at the time of developing the app, to have a better understanding of what data will be collected and hence what requirements are necessary to fulfill in order to develop apps that comply with GDPR. We remind the reader that the GDPR imposes obligations for protecting personal data. It requires more strict obligations (e.g., security mechanisms) when sensitive personal data (e.g., health information) is collected compared to less or non sensitive data, such as the user name invented by an individual for using a specific mobile app.
In response to RQ1, we found that the literature mostly focuses on high-level concepts of the model and mostly three concepts: (1) PERSONAL DATA ORIGIN, especially the collection of personal data; (2) RECIPIENTS, which has to do with the sharing of personal data; (3) LEGAL BASIS, and more specifically L2 CONSENT, which deals with the informed consent provided by the data subject. In contrast, several L2 concepts, such as WITHDRAW CONSENT or all the DATA SUBJECT RIGHTS, which are essential concerns of the GDPR regarding the protection of the data subjects, are not enough represented.
@@
RQ2 investigates the main research problems related to GDPR compliance addressed by the SE community. The main objectives observed in our work can be summarized across eight themes (T1–T8), further discussed in the rest of the section:
  1. T1.
    Tracing personal data in mobile apps
  2. T2.
    Analyzing app permissions
  3. T3.
    Analyzing the security and privacy of minors
  4. T4.
    Analyzing the implementation of informed consent
  5. T5.
    Analyzing users’ awareness on security and privacy
  6. T6.
    Analyzing the consistency between mobile apps and their privacy policies
  7. T7.
    Analyzing the implementation of the data minimization principle
  8. T8.
    Summarizing privacy policies
Table 7 lists the primary studies corresponding to each theme. While our analysis here focuses primarily on the objectives of the primary studies, we discuss below also the relevant concepts under each theme to better connect with our findings in RQ1. We identified relevant concepts using a deductive analysis. Recall from Section 4.4 that we adopted an integrated coding approach. For this RQ, we first synthesized the themes using inductive analysis, then we applied deductive analysis to map themes to the pre-defined concepts derived from the reference model.
Table 7: Overview of Themes Observed in the Literature (RQ2)
Theme Total Studies
T1 8 [1, 28, 41, 46, 50, 51, 52, 59]
T2 6 [2, 3, 20, 25, 27, 57]
T3 3 [9, 24, 32]
T4 11 [5, 7, 11, 21, 39, 42, 44, 45, 48, 49, 56]
T5 10 [16, 19, 22, 23, 26, 31, 33, 34, 36, 58]
T6 7 [18, 29, 35, 37, 38, 55, 60]
T7 7 [4, 8, 10, 12, 13, 14, 43]
T8 8 [6, 15, 17, 30, 40, 47, 53, 54]
5.2.1 T1. Tracing personal data in mobile apps
Primary studies often trace certain personal data to identify potential data leakage, i.e., unintended or unauthorized data exposure beyond the mobile app, such as its sharing with third parties. This leakage may lead to unlawful disclosure and/or data breaches.. This theme was observed in eight studies. Identifying which personal data to trace is a relevant research topic.
@@
Fostering this research theme is essential for understanding whether the mobile app is GDPR compliant when collecting personal data from the user. The analysis of source or byte code of mobile apps should be better aligned with the GDPR provisions as well as what has been stated in their respective privacy policies.
@@
We believe that this theme crosses multiple concepts, including PERSONAL DATA CATEGORY, PD ORIGIN, and CONSENT. Sharing data with third-party libraries is a closely related theme; in our reference model, this theme concerns the concept of RECIPIENT. In brief, the GDPR requires all recipients to be listed in the privacy policies of mobile apps. If the app is sharing personal data with third-party libraries, then this has to be fully transparent.
5.2.2 T2. Analyzing app permissions
@@
This theme investigates the access to personal data granted to the mobile app through permission requests. As mobile apps are becoming more pervasive, this theme is important for understanding what personal information the mobile app can access or collect indirectly through permission requests, e.g., to access information in other apps, risking that apps may gain access to data they were originally not allowed to access. Typical example is when the user ought to share location information to be able to use various apps such as navigation or fitness apps. This theme is related to the following concepts: PERSONAL DATA SECURITY, RECIPIENTS and PD ORIGIN, the latter two being among the most frequent concepts as per RQ1. We observed this theme in six primary studies.
5.2.3 T3. Analyzing the security and privacy of minors
@@
This theme focuses on minors, with an emphasis on the threats, risks, and challenges pertinent to this user group (e.g., obtaining valid informed consent), and the need to ensure their rights are adequately protected. This theme maps directly to the CHILDREN concept. We observed this theme in three primary studies.
5.2.4 T4. Analyzing the implementation of informed consent
@@
Consent is a well-studied concept. The GDPR emphasizes on collecting explicit consent from individuals, enabling them to make informed decisions with regard to the handling of their personal data. However, according to the reference model we apply in this SLR, CONSENT is only one of the several possible legal bases for data collection and processing (occurring in 32 studies).
@@
CONSENT is also the easiest to verify in software as it should be translated into interaction between the mobile app and the user. To investigate whether the practices in the mobile app code are GDPR compliant with respect to some other legal bases, a more comprehensive analysis of the app context must be conducted; such an analysis would not be necessarily limited to, e.g., the app’s source code. We note that, despite not being popular in the SE literature, legal bases other than consent (which are mentioned in only one paper out of 32) can play an important role in mobile apps development decisions.
5.2.5 T5. Analyzing users’ awareness on security and privacy
@@
This theme (observed in ten studies) surveys the users of mobile apps to study how well individuals are aware of their rights as well as common practices concerning privacy and security issues, often perceived in mobile apps’ behaviors. While this theme highlights the importance of individuals’ awareness about data protection, it cannot be directly mapped to our reference model.
@@
5.2.6 T6. Analyzing the consistency between mobile apps and their privacy policies
@@
Some papers investigate the consistency between the content of the privacy policy associated with an app against the actual behavior of that app. However, this research theme is often scoped to a use case such as the collection and sharing of sensitive personal data with third parties. Seven studies deal with this theme.
@@
Advancing this theme requires in-depth analysis of at least the following concepts: PERSONAL DATA CATEGORY, LEGAL BASIS, RECIPIENTS, PD ORIGIN, and DATA SUBJECT RIGHT. These concepts are related to the understanding of what personal data is collected, the source where it is obtained from, with whom it will be shared, what rights individuals can have on their personal data, and how these rights are implemented.
5.2.7 T7. Analyzing the implementation of the data minimization principle
Data minimization is an important guiding principle in GDPR: according to this principle, an app should not collect more personal data than necessary for its use. We observed this theme in seven primary studies.
@@
To ensure that no personal data is unnecessarily collected, the findings related to these two concepts should then be cross-checked against PERSONAL DATA CATEGORY, indicating the personal data types collected, and PD ORIGIN, indicating the source where the data is collected from.
@@
This theme (observed in eight studies) focuses on making privacy policies more accessible to individuals (e.g., end-users or engineers) by providing informative summaries or categorization of the policies.
While the primary studies under this theme cannot be one-to-one mapped to concepts in our reference model, we believe that the reference model can be useful to advance this research theme as it provides a structured hierarchy of the necessary details that must be identified in privacy policies.
In brief, this study aims to provide a comprehensive overview of the literature landscape addressing GDPR privacy concerns in mobile app research. While RQ1 serves as a starting point by mapping existing work to the concepts defined in our reference model, RQ2 examines the main objectives pursued by the primary studies. The rationale for linking RQ2 back to RQ1 is to offer a deeper insight into why certain concepts were addressed and others were not, driven by the overarching objectives of the primary studies.
In response to RQ2, we identified eight main themes in the literature, including tracing personal data for detecting potential data leakage, analyzing consent implementation and data minimization principles, checking the consistency between granted permissions, data use, and privacy policies, as well as topics addressing the privacy of minors, users’ privacy awareness, and the summarization of privacy policies.
@@
RQ3 explores the various contributions identified in the literature concerning GDPR. We distinguish in this SLR between technical and non-technical contributions: the former refers to automated or semi-automated approaches proposed to address specific GDPR-related privacy concerns, whereas the latter focus on sharing insights or recommendations, often derived from surveys, with the goal of advancing the research knowledge in this domain.
@@
Out of these studies, 34 are qualitative, relying on manual means for analyzing certain privacy concerns. However, some studies used simple tool support such as keyword-based or string search for extracting indicator terms, or applied third-party tools such as Mobile Security Framework (MobSF) [^11] https://github.com/MobSF/Mobile-Security-Framework-MobSF, FlowDroid [^12] https://github.com/secure-software-engineering/FlowDroid for static or dynamic code analysis, or Wireshark [^13] https://www.wireshark.org/ for network traffic analysis. Such an automation was strictly used to support their empirical studies or surveys.
As mentioned previously, most primary studies with non-technical contribution often described user studies or app surveys. The number of participants involved in user studies performed under this category vary between eight to 6,124 participants. Similarly, the number of apps considered in apps surveys vary, ranging from small app datasets (27 in [3], or 28 in [2]) to more than 10,000 apps in large-scale studies such as [12].
@@
The remaining 25 primary studies make technical contributions. They proposed automated or semi-automated approaches, regardless of whether they have resulted in full-fledged automated tools or prototypes. These approaches leverage different enabling technologies: model-driven engineering was used to enhancing business process modeling notation with privacy considerations [6], static and dynamic program analysis were used for detecting violations or vulnerabilities in mobile code [35, 41, 46, 51, 59], NLP was employed for analyzing privacy policies and app reviews [36, 44, 53, 57], machine learning (ML) and deep learning (DL) were used for classifying privacy policy provisions [30] and taint analysis for detecting potential data leaks in app code [1]. 13 of the primary studies with technical contributions leveraged a combination of these enabling technologies, particularly NLP, ML, DL, and static or dynamic program analysis. These studies pursued the following objectives:
  • detecting privacy issues in app reviews [16],
  • identifying data recipients [8],
  • detecting sensitive information disclosure in mobile apps [20],